OpenStudy (anonymous):

R-2k=15 solve of k Please can someone explain to me please

jimthompson5910 (jim_thompson5910):

the idea is to isolate k, so you need to move everything over to get k by itself any ideas how to start?

OpenStudy (isaiah.feynman):

k=15-R/2

OpenStudy (luigi0210):

@Isaiah.Feynman No giving answers

OpenStudy (luigi0210):

and @Reana \[\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile } \]

OpenStudy (isaiah.feynman):

@Luigi0210 Sorry sir!